The city of Waterbourne has made it a strategic goal to enhance livability in the city. The city has decided to undertake a project that will occur in two phases. During the first phase, the city will connect several of the parks with walking paths. During the second phase, the city will add water features or splash fountains to several of the parks.

The initial cost of the project will be $381,944. Annual utilities for maintaining the walking paths and keeping them lit during the early evening are estimated to be $13,458, with costs increasing by $214 each subsequent year. The paths will require repaving in year 9 at a cost of $55,214 and again in year 19 at a cost of $74,689.

The water features and splash fountains will be added to the parks at the end of year 9 at a cost of $294,551. The water utilities are estimated to be $18,544 per year, increasing by 13% each subsequent year. (The first payment for the water utilities will occur at the end of year 9.)

The city has secured a sponsorship from a local business to help pay for part of the project. The city will receive a grant from the business of $137,332 per year, increasing by 2% each subsequent year. The city will receive the grant for 7 years. (The city will receive the first grant payment at the end of year 1.)

Using a nominal annual interest rate of 2% compounded annually and a lifespan of 28 years, what is the present worth of the entire project?
Answer #1

Using Excel

Years Cash Flow
Investment utilities for path Water utilities Sponsorship Net cash flow
0 -381944.00 -3,81,944
1 -13458 137332 123874
2 -13672 140079 126407
3 -13886 142880 128994
4 -14100 145738 131638
5 -14314 148653 134339
6 -14528 151626 137098
7 -14742 154658 139916
8 -14956 -14956
9 -349765 -15170 -18544 -383479
10 -15384 -20955 -36339
11 -15598 -23679 -39277
12 -15812 -26757 -42569
13 -16026 -30236 -46262
14 -16240 -34166 -50406
15 -16454 -38608 -55062
16 -16668 -43627 -60295
17 -16882 -49298 -66180
18 -17096 -55707 -72803
19 -74689 -17310 -62949 -154948
20 -17524 -71132 -88656
21 -17738 -80379 -98117
22 -17952 -90829 -108781
23 -18166 -102636 -120802
24 -18380 -115979 -134359
25 -18594 -131056 -149650
26 -18808 -148094 -166902
27 -19022 -167346 -186368
28 -19236 -189101 -208337
NPW -10,97,324

Showing formula in Excel

Years Cash Flow
Investment utilities for path Water utilities Sponsorship Net cash flow
0 -381944 =B3+C3+D3+E3
1 -13458 137332 =B4+C4+D4+E4
2 =C4-214 =E4*1.02 =B5+C5+D5+E5
3 =C5-214 =E5*1.02 =B6+C6+D6+E6
4 =C6-214 =E6*1.02 =B7+C7+D7+E7
5 =C7-214 =E7*1.02 =B8+C8+D8+E8
6 =C8-214 =E8*1.02 =B9+C9+D9+E9
7 =C9-214 =E9*1.02 =B10+C10+D10+E10
8 =C10-214 =B11+C11+D11+E11
9 =-294551-55214 =C11-214 -18544 =B12+C12+D12+E12
10 =C12-214 =D12*1.13 =B13+C13+D13+E13
11 =C13-214 =D13*1.13 =B14+C14+D14+E14
12 =C14-214 =D14*1.13 =B15+C15+D15+E15
13 =C15-214 =D15*1.13 =B16+C16+D16+E16
14 =C16-214 =D16*1.13 =B17+C17+D17+E17
15 =C17-214 =D17*1.13 =B18+C18+D18+E18
16 =C18-214 =D18*1.13 =B19+C19+D19+E19
17 =C19-214 =D19*1.13 =B20+C20+D20+E20
18 =C20-214 =D20*1.13 =B21+C21+D21+E21
19 -74689 =C21-214 =D21*1.13 =B22+C22+D22+E22
20 =C22-214 =D22*1.13 =B23+C23+D23+E23
21 =C23-214 =D23*1.13 =B24+C24+D24+E24
22 =C24-214 =D24*1.13 =B25+C25+D25+E25
23 =C25-214 =D25*1.13 =B26+C26+D26+E26
24 =C26-214 =D26*1.13 =B27+C27+D27+E27
25 =C27-214 =D27*1.13 =B28+C28+D28+E28
26 =C28-214 =D28*1.13 =B29+C29+D29+E29
27 =C29-214 =D29*1.13 =B30+C30+D30+E30
28 =C30-214 =D30*1.13 =B31+C31+D31+E31
NPW =NPV(2%,F4:F31)+F3